OpenStudy (anonymous):

How many grams of calcium phosphate can be produced when 78.5 grams of calcium hydroxide reacts with excess phosphoric acid? Unbalanced equation: H3PO4 + Ca(OH)2 ---> H2O + Ca3(PO4)2 @DanJS

OpenStudy (danjs):

try balancing the equation, i will brb in 2 min, have to change over laundry

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\(\sf H_3PO_4 + 3Ca(OH)_2 \rightarrow H_2O + Ca_3(PO_4)_2\) Hmmm I am getting stuck balancing this

OpenStudy (danjs):

Do the cation metals first

OpenStudy (danjs):

then groups likeOH and PO4

OpenStudy (danjs):

then finally oxygen and hydrogen

OpenStudy (danjs):

2,3,6,1 try that

OpenStudy (anonymous):

\(\sf \sf 2H_3PO_4 + 3Ca(OH)_2 \rightarrow 6H_2O + Ca_3(PO_4)_2\)

OpenStudy (danjs):

really no trick to doing that, just which atoms are more important to balance first....and you can use fractions if needed, then multiply through at the end to get all whole numbers

OpenStudy (anonymous):

It still isn't balanced though. There are 8 H on the left but 12 on the right.

OpenStudy (danjs):

look again... (OH)2 has 2 H each, and you have 3 molecules, that is 6 H atoms there

OpenStudy (danjs):

H: 6 + 6 = 12

OpenStudy (anonymous):

Oh ok

OpenStudy (danjs):

Just have to keep tryin at those, adding coefficients and changing them around until it checks out

OpenStudy (danjs):

doing, metal cations first, then groups(OH) and (PO4) , then oxygen, then hydrogen

OpenStudy (danjs):

So now you have the balanced equation , we can do the stoichiometry

OpenStudy (danjs):

2H3PO4 + 3Ca(OH)2 → 6 H2O + Ca3(PO4)2 78.5g Ca(OH)2 Excess = infinite H3PO4

OpenStudy (danjs):

Calculate the Molar Masses of each compound in question

OpenStudy (danjs):

and i will start drawing up the solution

OpenStudy (anonymous):

H3PO4 = 98 g/mol Ca(OH)2 = 74.093 g/mol H2O = 18.01528 g/mol Ca3(PO4)2 = 310.1767 g/mol

OpenStudy (danjs):



OpenStudy (danjs):

I used your molar masses, assuming they are right.. :)

OpenStudy (anonymous):

109.54

OpenStudy (danjs):

general format: (Grams reactant) x (Molar Mass) x (mole Ratio) x ( Molar mass)

OpenStudy (danjs):

yep that is what i got

OpenStudy (danjs):

grams of calcium phosphate

OpenStudy (danjs):

see how the units cancel, and you are left with what you are looking for, grams of calcium phosphate?

OpenStudy (anonymous):

Yes

OpenStudy (danjs):

That is all you have to do.. Start by writing what they give you... then multiply by fractions of 1 and end with the unit they are lookig for

OpenStudy (danjs):

if you want 'how many molecules' use as the last fraction this instead \[\frac{ 6.022 x 10^{23} ~molecules ~~Ca(PO4)2 }{ 1 mol ~~Ca(PO4)2 }\]